The population of a city increases by 4,000 people each year. In 2025 the population is projected to be 450,000 people,. What is the equation that gives the city's population p ( in thousands of people) x years after 2010?
Answer by TimothyLamb(4379)

population equation (in thousands of people), starting in 2025:
p = 4x + 450
to use this equation, x = (year - 2025)
---
find population (in thousands of people), in 2010:
p(2010) = 4*(2010 - 2025) + 450
p(2010) = 4*(-15) + 450
p(2010) = 390
---
answer:
---
population equation (in thousands of people), starting in 2010:
p = 4x + 390
to use this equation, x = (year - 2010)
---
check the answer:
by calculating the population (in thousands of people), in the year 2025:
p = 4x + 390
x = 2025
p = 4*(2025 - 2010) + 390
p = 4*(15) + 390
p = 450
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php